Circular No. 8577

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2003 EL_{61}, 2003 UB_{313}, AND 2005 FY_{9}

Three relatively bright transneptunian objects have been reported by M. E. Brown, C. A. Trujillo, and D. Rabinowitz using the Palomar 1.2-m Schmidt telescope (+ Quest camera). The object 2003 UB_{313}, currently at heliocentric distance r = 97 AU and red mag 18.5, has absolute magnitude H = −1.1 (Pluto has H = −1.0); observations, the following orbital elements (P = 557 yr), and an ephemeris appear on MPEC 2005-O41.

Epoch = 2005 Aug. 18.0 TT

T = 2257 Jan. 26.1837 TT  ω = 151.^{o}3115
e = 0.441613  Ω = 35.8750 }2000.0
q = 37.808 AU  i = 44.1770

MPEC 2005-O42 contains observations, orbital elements (T = Sept. 2130, q = 38.7 AU, ω = 246^{o}, Ω = 79^{o}, i = 29^{o}, e = 0.15, P = 308 yr), and an ephemeris for 2005 FY_{9} (which is currently at red mag 17).

MPEC 2005-O36 contains astrometric observations for an object, designated 2003 EL_{61}, of red mag 17 first reported by P. Santos-Sanz, J. L. Ortiz, and F. J. Aceituno via three CCD frames taken in March 2003 with a 0.36-m Schmidt-Cassegrain reflector at the Sierra Nevada Observatory in Granada, Spain, with additional observations identified from images obtained at several different sites back to 1955. MPEC 2005-O36 also contains orbital elements (T = Dec. 2133, q = 35.2 AU, ω = 240^{o}, Ω = 122^{o}, i = 28^{o}, e = 0.19, P = 285 yr) and an ephemeris for 2003 EL_{61}. This object was also found by Brown et al. at Palomar on 2004 May 6 [Rabinowitz et al. 2005, Bull. A.A.S. 37(3), Abstr. 56.12], who also obtained observations at Keck Observatory during Jan.-June 2005 of a satellite with orbital period 49.1 days and semimajor axis ≈ 49500 km, which indicate a total system mass ≈ 29 percent that of the mass of the Pluto-Charon system [Bouchez et al. 2005, Bull. A.A.S. 37(3), Abstr. 56.02; Brown et al. 2005, Ap.J. Let., submitted].

COMET C/2005 N6 (SOHO) Another very faint and diffuse tailless Kreutz sungrazing comet has

been found on SOHO images (cf. IAUC 8576; TC = T. Chen).

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 N6 July 6.354 6^{h}54.^{m}1 +21^{o}22′ C2 TC 2005-O26
2005 July 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green